OpenStudy (anonymous):

write the equation of the line with m = 9, and y-intercept (0,-6) in slope-intercept form. A. y = 9x - 6 B. y = 6x - 9 C. y = -9x + 6 D. y = -6x + 9

OpenStudy (hoblos):

y=mx +b y = 9x - 6

sam (.sam.):

\[y-y _{1}=m(x-x _{1})\] \[y-(-6)=9(x-0)\] \[y=9x-6\]

Directrix (directrix):

A. y = 9x - 6 y = mx + b